OpenStudy (anonymous):

PLEASE HELP - What is the equation for sound volume in decibels? is it dB = 20log(p2/p1) or dB = 10log(p2/p1) ?? does the volume of sound count as sound pressure?

OpenStudy (vincent-lyon.fr):

with P being a power/m², use factor 10 (real definition) with p being the amplitude of the acoustic pressure, use a factor 20. because power P is a quadratic quantity, i.e. proportional to p² \(10\;lg\Large\left(\frac{p_2}{p_1}\right)^2\)= \(20\;lg\Large\left(\frac{p_2}{p_1}\right)\)